OpenStudy (anonymous):

When Θ = 5 pi over 6, what are the reference angle and the sign values for sine, cosine, and tangent? Θ' = negative pi over 6; sine and cosine are positive, tangent is negative. Θ' = 5 pi over 6; sine and tangent are positive, cosine is negative Θ' = pi over 6; sine is positive, cosine and tangent are negative Θ' = negative 5 pi over 6; sine is positive, cosine and tangent are negative @satellite73 plz help

OpenStudy (anonymous):

i think it's C but not sure

OpenStudy (anonymous):

i never know about reference angles but i bet we can figure it out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have to look it up i think it may be C too

OpenStudy (anonymous):

yeah go with C

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

reference angles the stupidest damn thing no use at all

OpenStudy (anonymous):

haha how i feel about math all the time xD

OpenStudy (anonymous):

yeah even on christmas eve! don't you ever take a break?

OpenStudy (anonymous):

i dont celebrate it lol

OpenStudy (anonymous):

do u know the answer?

OpenStudy (anonymous):

yeah it is C

myininaya (myininaya):

wow i thought I was the only one who didn't use reference angles

OpenStudy (anonymous):

it is ridiculous

OpenStudy (jhannybean):

I use reference angles...

OpenStudy (anonymous):

spend more time figuring out the "reference" when you can just use the unit circle

OpenStudy (jhannybean):

Well, then again I already know what they are haha.

OpenStudy (anonymous):

just curious when you would use it would you use it to find \(\cos(\frac{5\pi}{6})\) and if so, how?

myininaya (myininaya):

\[\cos(\frac{5\pi}{6})=-\cos(\frac{\pi}{6})\]

OpenStudy (anonymous):

too many rules, just look at the unit circle or do it menally

OpenStudy (jhannybean):

I just refer back to Q1 to find \(\cos\left(\frac{\pi}{6}\right)\) to either find my angle measure or the cosine or sine value correlated with it, just to cross check.

OpenStudy (jhannybean):

I've gotten so used to referring back to Q1 to crosscheck my values I guess it's become a habit.

OpenStudy (anonymous):

it always scares me when someones picture changes in the middle of a thread

OpenStudy (jhannybean):

Haha, It was @Zale101's gift to me this year, had to use it.

OpenStudy (anonymous):

The stupid way that we were taught about reference angles for this was when you have something such as 5pi/6 you could find the reference angle by just doing this